OpenStudy (anonymous):

apply z transform to x[n] = a^(n+1) u[n+1] step by step please

OpenStudy (anonymous):

u[n] is a unit step function it is used as a bounded input

OpenStudy (anonymous):

Is this a bilateral Z transform?

OpenStudy (anonymous):

it is bilaterial with out the unit step function

OpenStudy (anonymous):

the unit step is for a input bound

OpenStudy (anonymous):

the unit step changes the bounds

OpenStudy (anonymous):

I really don't know what I'm doing. I'll leave this to someone else.

OpenStudy (anonymous):

kk, ty for trying

OpenStudy (ash2326):

Let the discrete function be a(n) Z transform is defined as \[\Large H(z)=\Sigma_{n=-\infty}^{n=\infty} a(n)z^{-n}\] Do you get this?

OpenStudy (anonymous):

yes

OpenStudy (ash2326):

we have \[a[n]=a^{n+1} U(n+1)\] so the discrete function is defined from \(n=-1\ to\ n=\infty\)

OpenStudy (anonymous):

ok, i understand there is a time shift n+1 = 0 n=-1

OpenStudy (ash2326):

yeah:)

OpenStudy (ash2326):

Let's plugin the values

OpenStudy (anonymous):

k

OpenStudy (ash2326):

\[\Large H(z)=\Sigma_{n=-1}^{n=\infty} a^{n+1}z^{-n}\] \[\Large H(z)=a\times \Sigma_{n=-1}^{n=\infty} a^{n}z^{-n}\] or \[\Large H(z)=a\times \Sigma_{n=-1}^{n=\infty} (az^{-1})^n\] do you get till here?@mathman12

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

i got to here and try to match a infinite geometric series and got something strange

OpenStudy (ash2326):

if \[|a|\le 1\] then this z transform will exist, other wise it won't exist

OpenStudy (anonymous):

i tried matching \[\sum_{k=0}^{oo} x^k = 1/(1-x)\]

OpenStudy (ash2326):

you can do that [\[\Large H(z)=a\times ((az^{-1})^{-1}+\Sigma_{n=0}^{n=\infty} (az^{-1})^n)\] now we get [\[\Large H(z)=a\times ((\frac z a+\Sigma_{n=0}^{n=\infty} (az^{-1})^n)\] Now use the formula for the last part

OpenStudy (ash2326):

I took out n=-1 from the sum and then the sum becam from n=0 to n= (\infty\)

OpenStudy (ash2326):

n=\(\infty\)

OpenStudy (anonymous):

can't you subtract it from a^(n+1)z^(-n) as a^(n+1-1)z^(-(n-1)) ?

OpenStudy (anonymous):

to compensate for the adding to the bound you subtract from the exp?

OpenStudy (ash2326):

let me think

OpenStudy (ash2326):

that won't work

OpenStudy (ash2326):

Read my post , if you don't understand. I'll help you:)